Slope-1/2, ordered pair (2,3)
Answer by Cromlix(4381)

Hi there,
y - 3 =-1/2(x - 2)
y - 3 = -1/2x + 1
y = -1/2 + 1 + 3
y = -1/2x + 4
or
2y = -x + 8
Hope this helps :-)